OpenStudy (anonymous):

How many liters of water must evaporate from 50 liters of an 8% salt solution to make 25% salt solution?

OpenStudy (anonymous):

8% of 50= 4. means we have 4 parts of salt in 50 litres of water. since salt content will remain same. so 25% of x= 4. so answer ths equation gives is 16. therefore, 34 litres of water should evaporate.

OpenStudy (anonymous):

mass of salt in 50 liters of 8% = 50*80 = 4000 for a 25% solution volume woold be 4*4000 = 16000 so 50 -16 = 34 will evaporate